OpenStudy (anonymous):

Medal/Fan rewarded for the person who can help me with this question. Your friend runs up to you, scared that he is not ready for the upcoming quadratics test. To help him study,you will create four different quadratic functions. Then demonstrate how to rewrite each function as a group of factors, if possible. a. The function f(x) is a difference of squares b. The function g(x) is a sum of squares c. The function h(x) is a perfect square trinomial d. The function j(x) can only have a GCF factored out of it.

OpenStudy (anonymous):

f(x) = x^2 – 4 = (x – 2)(x + 2) g(x) = y^2 + 5^2 = y^2 + 25 h(x) = (x-5)x(x-5) = x^2 -10x + 25 j(x) = 15x+30 = 15(x+2)

OpenStudy (anonymous):

Thank you so much :)

OpenStudy (anonymous):

No problem :)